OpenStudy (anonymous):

How does one calculate a ||f1||??? doing the Gram-Schmidt Algorithm

OpenStudy (amistre64):

gram schmidt eh .... sounds orthogonal to me

OpenStudy (amistre64):

if you are doing an algorithm, isnt that just a matter of following the instructions?

OpenStudy (anonymous):

What do you mean? Like I have the equation and all, I just don't know how my prof calculated ||f1||

OpenStudy (anonymous):

From what I can see you just multiple f1 by itself

OpenStudy (amistre64):

can you post the problem, to give this a little more context? i know || is a magnitude of a vector but i see no convention of f1

OpenStudy (amistre64):

||f1|| would be the square root of : f1 dot f1

OpenStudy (anonymous):

Oh lol sorry. FInd the QR factorization of A = \[\left[\begin{matrix}1 & 3&2 \\ -1&1&0\\0&2&1\\1&1&4\end{matrix}\right]\]

OpenStudy (anonymous):

so I followed along an example in my notes and did out to find f1, f2 and f3 with the equations f1=v1 f2=\[f1\times f1\left( v2 - \frac{ v2 \times f1 }{ f1 \times f1 } \times f1 \right)\]

OpenStudy (anonymous):

and f3=\[f1 \times f2 \left( v3-\frac{ v3 \times f1 }{ f1 \times f1 } \times f1 - \frac{ v3 \times f2 }{ f2 \times f2 } \times f2\right)\]

OpenStudy (amistre64):

the first vector is generally taken to be the first column vector \[u_1 =a_1 = [1~-1~0~ 1]\] \[e_1 =\frac{a_1}{|a_1|}\text{ the unit vector of }a_1\] \[u_2 =a_2-(a_2\cdot e_1)e_1~~...\]

OpenStudy (anonymous):

so I solved for those three... then in my notes I seen that I need to find q1, q2 and q3 by q1 = \[\frac{ f1 }{ ||f1|| }\] and q2 =\[\frac{ f2 }{ ||f2|| }\] and q3= \[\frac{ f3 }{ ||f3|| }\]

OpenStudy (anonymous):

But I'm not sure what ||f1|| means

OpenStudy (anonymous):

Theses are all vectors btw lol

OpenStudy (amistre64):

||v|| is notation for the length of a vector \(\vec v\). any vector, that is divided by its own length, become a unit vector of itself. you can use it as a blank slate to scale as needed

OpenStudy (anonymous):

How do you like solve it though? Do you just times the vector by itself?

OpenStudy (amistre64):

if youve already determined the vectors for the fs, make them into unit vector you can acheive the length of a vector by dotting it to itself and taking the sqrt of the results for example: let v = (3,1,2) v.v = 3,1,2 3,1,2 ------ 9+1+4 = 14 |v| = sqrt(14)

OpenStudy (amistre64):

square add sqrt is also a good memory devies

OpenStudy (anonymous):

Ohhhhh... So for example if f1=(1,-1,0,1) it'd be sqrt(3) ?

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

Oh okay thank you :)

OpenStudy (amistre64):

youre welcome, good luck :)